Need Help!!! use the discriminant to determine whether the following equations have solutions that are: two different rational solutions; two different irrational solutions; exactly one rational solution; or two different imaginary solutions.
s^2-5s-6=0

Answer by stanbon(75887)

use the discriminant to determine whether the following equations have solutions that are: two different rational solutions; two different irrational solutions; exactly one rational solution; or two different imaginary solutions.
s^2-5s-6=0
----
a = 1 ; b = -5 ; c = -6
--------
b^2-4ac = 25 - 4*1*-6 = 49
---
Since 49 is positive there are two different Real solutions.
---
The discriminant does not quarantee the solutions
will be rational, only that they will be Real.
================================================
Cheers,
Stan H.
=================